OpenStudy (anonymous):

A number is divided by 3. If 8 is subtracted from the quotient you get -6 A. -6 B. 3 C. 6 D. 42

OpenStudy (anonymous):

Consider the number as x. x/3-8=-6 x/3=-6+8 x/3=2 x/3*3=2*3 x=6 C is the right answer.

OpenStudy (texaschic101):

x/3 - 8 = -6 (multiply the equation by 3) x - 24 = -18 (add 24 to both sides) x = -18 + 24 x = 6